# Causality, Emergence, Causal Emergence, Causal Emergence Identification, Causal Representation Learning, World Model

### 28 Dec 2023 v1

@@ 2.2 Emergence
Emergence, known as "the whole is greater than the sum of parts" Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro and macro levels of organization. Therefore, the scale of a system (micro or macro) and the cross-level interactions must be considered when we discuss emergence. This may increase the complexity of study for emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al. (2019); Ross et al. (1908); Holland (1996)
@@
Emergence has deep connections with causation Fromm (2005); Bedau (1997); Bedau and Humphreys (2008); Bedau (2002). First, emergence emphasizes the new properties arising from the interactions between the components in the system, rather than the components themselves. Consequently, all emergent properties and processes at the macro-level can, in principle, be comprehended as the causal effects resulting from interactions between individuals at the micro-level. This is what the philosophical notion “supervenience” tries to describe.
@@
However, it does not mean that all the emergent phenomena on macro-level can be easily attributed to micro-level individuals and their interactions. Two major reasons can account for this. The first reason is due to the principle of computational irreducibility proposed by Wolfram Wolfram et al. (2002) which claims that although we know the emergent properties are out of a set of simple rules, we can not find a short-cut way in principle to predict the results except to implement the rules. For example, complex behaviors (eg. herding and aligning) of the Boid model are the results of simple rules, but these behaviors can not be reduced to individual behaviors in a simple way except to simulate it Reynolds (1987).
@@
The second reason is that emergent behaviors may be determined by other emergent properties on macro-level Hoel et al. (2013); Rosas et al. (2020). For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because the severe lack of fossil oil may cause inflation. Therefore, the cause of an emergent property (rice price) is another emergent macro-level coarse-grained variable and it means new causation on macro-level coarse-grained variables can be observed as pointed by the theory of causal emergence. And, interestingly, the causation in different levels must be consistent to explain the micro-level event of the price of rice. Otherwise, double causation Merton (1973) fallacy may appear. This consistency between different scale causality is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
@@
In order to better understand emergence, Bedau et al Bedau (1997, 2002) classify emergence into three categories according to the causal interactions between micro and macro, which are nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996) and strong emergence O’Connor (1994); Kim (1992).
@@
In all these categories, the notion of nominal emergence is least controversial, which can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994). For example, the patterns which are composed of many pixels on a screen are nominal emergence. We can say nominal emergent patterns as “supervenient” because all the macro-level properties are from individuals.
@@
While weak emergence refers to macro-level properties or processes that can be derived through interactions between individual components but in a complicated way such that they are hardly reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996). For the weak emergent patterns or processes, the causes may come from both micro-level and macro-level, therefore, emergent causation may coexist with micro-level causation.
@@
There are many debates on strong emergence, which refers to properties or processes on macro-level that can not be reduced to micro-level properties in principle, including the interactions of individuals O’Connor (1994); Kim (1992). Thus, the collective behaviors of boids are not strong emergent because they are the results of interactions between Boid model. This notion is controversial because it rejects any mechanistic explanations for strong emergent properties, which can not be explained by micro-level variables. Second it raises some long standing philosophical debates about “causal fundamentalism”  Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
@@
Jochen Fromm further explains strong emergence as the causal effect of downward causation Fromm (2005). Consider a system contains three-different scales: micro, meso, and macro. Downward causation refers to the causal power from the macro-level to the messo-level or from the messo-level to the micro-level. Thereafter, although the strong emergent properties or processes on the meso-level are supervenient to micro-level properties and interactions, their causes are from the macro-level and therefore a second order patterns or processes. However, there are many debates on the notion of downward causation itself, e.g.  Yurchenko (2023); Bedau and Humphreys (2008).
Figure 2: Various types of causation that exist within multi-scale complex systems. In this figure, solid arrows represent common causation, which is generally accepted without much controversy. Dotted arrowed lines indicate a form of “causation” driven by supervenience, while dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
@@
1. Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
2. Intra-level causation: This refers to causal effects occurring within the same scale or level.
3. Downward causation: Here, macro-level properties have influences on micro-level properties.
In Figure 2, the black solid lines represent causation without controversy. However, according to a recent paper by Yurchenko  Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality, but rather reasons or explanations. Yurchenko introduces the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers. Instead, they serve as reasons or explanations from the perspective of an observer. While, Yurchenko accepts the intra-level causation in the meso- or macro-level as “real” as the ones in micro-level. He even coined the term “causal equivalent principle” to represent the idea that all the intra-level causation except cross-leveled ones should be accepted and are equivalent each other.
However, other scholars think that the emergent causation, the intra-level causation in the meso- or macro- level, depends on observers Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022). There are much controversy on emergent cross-leveled causation.
@@ 3 Quantifying Emergence by Causality
Since causality and emergence have a strong connection, and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence. In this section, we will review several frameworks to quantify multi-scale causality and emergence.
@@ 3.1 Early Related Works
Before the theory of causal emergence was proposed by Hoel et al, there were some related works that have introduced very similar ideas as causal emergence theory. For example, Crutchfield et al’s computational mechanics theory considered causal states which are the partitions of the state space and may have good predictions. While, Seth et al proposed G-causality to quantify emergence by using Granger’s causality. We will introduce them in detail.
@@ 3.2 Computational Mechanics
The theory of computational mechanics proposed by Crutchfield, Shalizi, and Feldman et al. tries to formulate this kind of emergent causation in a quantative framework Crutchfield and Young (1989). In some sense, computational mechanics can be understood as the inverse of statistical mechanics because what statistical mechanics do is to derive macro-level consequences from micro-dynamics, and the inverse process done by computational mechanics is to construct a minimal causal model from the observations of a stochastic process that can generate the observed time series.
Let’s assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$. We can divide it into two segments beginning from time step $t$: the history before $t$ which is denoted as $\overleftarrow{s_{t}}$ and the future after $t$ which is denoted as $\overrightarrow{s_{t}}$. If the process is stationary, we can drop the denotation of $t$. Thus, all the possible histories $\overleftarrow{s}$ form a set which is denoted as $\overleftarrow{S}$ and all the futures $\overrightarrow{s}$ form a set $\overrightarrow{S}$.
We aim to establish a model that can reconstruct and predict observed random sequences, with a higher accuracy being desirable. However, the randomness of the sequences prevents us from achieving a perfect reconstruction unless we record every randomly occurring character. This would make the model excessively long. To preserve useful information as concisely as possible, we need a coarse-grained mapping that captures the ordered structure in the random sequences, known as patterns Crutchfield (1994). We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets which form a set $\mathcal{R}$. Any subset in $R\in\mathcal{R}$ is a state. We define a function from histories to states, $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$. Thus $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
@@
What kind of state set can achieve the best balance between predictiveness and parsimony? We can introduce an equivalent relationship which is called causal equivalence Crutchfield (1994). Concretely, we say $\overleftarrow{s}$ and $\overleftarrow{s}^{\prime}$ are causally equivalent if and only if:
@@
This equivalent relationship can partition all the histories into equivalent classes, and they are defined as causal states. We denote all the causal states of a history $\overleftarrow{s}$ as: $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}}$ which is a subset of the histories.
For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that make the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
@@
where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and the notation $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$ in which the first symbol emitted is $s$111we use slightly different symbols compared with reference  Shalizi (2001). Hence, by combining: $\mathcal{S^{\prime}}=\mathcal{S}_{j}$ and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$, we ensure that all subsequent causal states will have the identical initial emitted symbol $s$. We denote the set $\{T^{(s)}_{ij}:s\in\mathcal{A}\}$ by $T$. And the definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$) and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
Up to now we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories. The remaining task is to show that it is in some sense optimal. It has the important properties of being minimally predictive, maximally statistically complex, and minimally stochastic Shalizi (2001).
We shall compare the causal state $\mathcal{S}$ with any state $R\in 2^{\overleftarrow{\mathcal{S}}}$, to show that for the purpose of predicting the future, the causal states does a better job, i.e. gives more information. This property can be formulated as a mathematical theorem which is called the maximal predictability theorem. This is stated as if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
@@
Where, $H(\cdot|\cdot)$ is the conditional entropy, $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future. Thus, the inequality shows that the uncertainty of $\overrightarrow{\mathcal{S}}^{L}$ given the causal state $\mathcal{S}$ is less than all other states.
After achieving optimal predictiveness, the causal state set remains the one with minimal statistical complexity. We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$ which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
@@
We next show that the causal states are minimally stochastic. That is to say, compared with other competitors with the same ability of predicting the future, the causal states and their transition dynamics, have the least uncertainty. Then we have minimal stochasticity theorem which is expressed as: for all prescient rivals $\hat{\mathcal{R}}$,
@@
Where $\mathcal{S}^{\prime}$ and $\hat{\mathcal{R}}^{\prime}$ are the next causal state of the process and the next state, respectively. This means causal state and $\epsilon$-machine gives the best intrinsic determinism.
Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from observed data? The authors introduced a hierarchical machine reconstruction algorithm in Crutchfield (1994), details are ignored.
Although this algorithm may not be applicable in all operational scenarios, the authors have presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
It is interesting to compare the theory of computational mechanics with causal emergence. Indeed, we can understand all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states. Thus the function $\eta$ that can map a history $\overleftarrow{s}$ to a state $R$ is a possible coarse-graining strategy.
It is worth to be pointed out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can has at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history. Therefore, $\epsilon$ is similar to the notion of effective coarse-graining strategy in reference  Zhang and Liu (2022) (see section 4.1) and the causal transitions $T$ are the corresponding effective macro-level dynamics. And the feature of minimal stochasticity characterize the deterministic property of the macro-dynamic. This is characterized by effective information (EI) in causal emergence theory.
Although the clear definition and quantitative theory of emergence are not given, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003). In Crutchfield (1994), the authors explain that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer. Additionally, they differentiate intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.

### 25 Feb 2024 v3

@@ 2.2 Emergence
Emergence, known as “the whole is greater than the sum of parts” Pepper (1926), is a central concept in many philosophical and theoretical discussions about the nature of complexity and the relationship between micro- and macro-levels of organization. Therefore, the scale of a system (micro or macro) and the cross-level interactions must be considered when we discuss emergence. This may increase the complexity of the study of emergence Winning and Bechtel (2019); Hendry (2019); Huxley (1947); Mill (2011); Gibb et al. (2019); Ross et al. (1908); Holland (1996).
Although many examples of emergence have been pointed out in various fields Holland (1996, 2000), including the herding behaviors of birds Reynolds (1987), collective behaviors of simple computer programs Gardner (1970), and emerging ability of large language models Wei et al. (2022), a unifying understanding of this phenomenon does not exist right now. People try to understand emergence from different aspects, including self-organization Luhmann (1995), order out of disorder Crutchfield (1994), and causality. In this review, we mainly focus on the last point of view.
@@
Emergence has deep connections with causation Fromm (2005); Bedau (1997); Bedau and Humphreys (2008); Bedau (2002). First, emergence emphasizes the new properties arising from the interactions between the components in the system, rather than the components themselves. Consequently, all emergent properties and processes at the macro-level can, in principle, be comprehended as the causal effects resulting from interactions between individuals at the micro-level. This is what the philosophical notion of “supervenience” tries to describe.
@@
However, this does not mean that all emergent phenomena at the macro-level can be easily attributed to micro-level individuals and their interactions. Two major reasons can account for this. The first reason is due to the principle of computational irreducibility proposed by Wolfram Wolfram et al. (2002), which claims that although we know emergent properties arise from a set of simple rules, there is no shortcut, in principle, to predict the results other than implementing the rules. For example, the complex behaviors (e.g., herding and aligning) of the Boid model are the results of simple rules, but these behaviors cannot be reduced to individual behaviors simply without simulating them Reynolds (1987).
@@
The second reason is that emergent behaviors may be determined by other emergent properties at the macro-level Hoel et al. (2013); Rosas et al. (2020). For example, although the price of rice in the market is determined by all the interactions (bargaining) between buyers and sellers, it is also the causal effect of the price of fossil oil because severe shortages of fossil oil can cause inflation. Therefore, the cause of one emergent property (rice price) is another emergent macro-level coarse-grained variable. This means new causation can be observed between macro-level coarse-grained variables, as pointed out by the theory of causal emergence. Interestingly, the causation at different levels must be consistent to explain the micro-level event of fluctuating rice prices. Otherwise, a double causation Merton (1973) fallacy may appear. This consistency between causality at different scales is called the causal equivalence principle in Yurchenko’s latest article Yurchenko (2023).
@@
In order to better understand emergence, Bedau et al. Bedau (1997, 2002) classified emergence into three categories according to the causal interactions between the micro- and macro-levels: nominal emergence Harré (1985); Baas (1994), weak emergence Bedau (1997); Newman (1996), and strong emergence O’Connor (1994); Kim (1992).
@@
Of these categories, the notion of nominal emergence is the least controversial. It can be described as a kind of property that can be possessed by macro-level patterns or processes but not by their micro-level components Harré (1985); Baas (1994). For example, pixel patterns on a screen are nominal emergent properties. We can consider such nominal emergent patterns as “supervenient” because all macro-level properties derive from individuals.
@@
Weak emergence refers to macro-level properties or processes that derive through interactions between individual components in a complicated way such that they cannot be easily reduced to micro-level properties due to the principle of computational irreducibility Bedau (1997); Newman (1996). For weak emergent patterns or processes, the causes may come from both the micro- and macro-levels; therefore, emergent causation may coexist with micro-level causation.
In the discourse surrounding emergence, weak emergence is generally more widely accepted than strong emergence. However, there are some concepts and descriptions that remain somewhat ambiguous Héder and Paksi . For instance, in Bedeau’s original definition, a property is considered weakly emergent if and only if it can be derived from micro-dynamics through simulation. Nevertheless, the term “simulation” itself can be interpreted in different ways—whether it refers to digital simulations or analog simulations is not explicitly implied. Bedeau further elaborated on this concept by stating that if a macro-level property can, in principle, be simulated without requiring an actual simulation, then it can be considered weakly emergent Bedau (1997). However, even this explanation is not entirely clear-cut.
@@
There are more debates on strong emergence, which refers to properties or processes at the macro-level that cannot, in principle, be reduced to micro-level properties, including the interactions of individuals O’Connor (1994); Kim (1992). Thus, Boid’s collective behaviors are not strongly emergent because they are the result of interactions within the Boid model. This notion is controversial because it rejects any mechanistic explanations for strong emergent properties, which cannot be explained by micro-level variables. Furthermore, it raises some long-standing philosophical debates about “causal fundamentalism” Jackson and Pettit (1992) and “supervenience” Kim (1999), and the existence of strong emergence remains an open problem due to the scarcity of concrete examples.
@@
Jochen Fromm further explained strong emergence as the causal effect of downward causation Fromm (2005). Consider a system that contains three different scales: micro, meso, and macro. Downward causation refers to the causal power from the macro-level to the meso-level or from the meso-level to the micro-level. Consequently, although the strong emergent properties or processes at the meso-level are supervenient to micro-level properties and interactions, their causes derive from the macro-level and, therefore, second-order patterns or processes. However, there are many debates on the notion of downward causation itself, e.g., Yurchenko (2023); Bedau and Humphreys (2008).
@@
  1. 1.
    Upward causation: This involves the supervenience relation, where macro-level effects can be attributed to micro-level variables.
  2. 2.
    Intra-level causation: This refers to causal effects occurring within the same scale or level.
  3. 3.
    Downward causation: Here, macro-level properties influence micro-level properties.
Figure 2: The various types of causation that exist within multi-scale complex systems. In this figure, solid arrows represent common causation, which is generally accepted without much controversy. Dotted arrowed lines indicate a form of “causation” driven by supervenience, whereas dashed arrowed lines represent emergent causation, occurring either at an intra-level or downward.
In Figure 2, the black solid lines represent causation without controversy. However, according to a recent paper by Yurchenko Yurchenko (2023), the causation represented by the dashed and dotted lines (except the intra-level ones in the meso-level) may not actually be causality but rather reasons or explanations. Yurchenko introduced the term “reason” to distinguish it from the conventional understanding of causality, as cross-level causal effects may not possess causal powers. Instead, they serve as reasons or explanations from the perspective of an observer. Yurchenko accepted intra-level causation at the meso- or macro-level as “real”, similar to causation at the micro-level. He even coined the term “causal equivalent principle” to represent the idea that all intra-level causations, except those that are cross-level, should be accepted and are equivalent to each other.
However, other scholars think that compared to the causal relationship at the micro-level, the intra-level causation at the meso- or macro-level is problematic because the coarse-graining strategy is relevant Bonabeau and Dessalles (2011); Dewhurst (2021); Eberhardt and Lee (2022). As demonstrated in the numeric example in Section 3.5.4, the causation measured by EI is dependent on the coarse-graining method. Hoel et al. tried to avoid this problem by introducing EI maximization. Indeed, EI maximization serves as an objective standard for selecting coarse-graining and macro-dynamics; however, a unique solution is required. We discuss this problem further in Section 5.2.
@@ 3 Quantifying Emergence by Causality
Since causality and emergence have a strong connection and causality has multiple quantitative frameworks and measures, it becomes natural to use causality to quantify emergence. In this section, we review several frameworks used to quantify multi-scale causality and emergence.
@@ 3.1 Early Related Works
Before the theory of causal emergence was proposed by Hoel et al., some works introduced very similar ideas to causal emergence theory. For example, Crutchfield et al.’s computational mechanics theory considered causal states, which are the partitions of the state space and may have good predictions. Furthermore, Seth et al. proposed G-causality to quantify emergence using Granger’s causality. We discuss them in detail below.
@@ 3.2 Computational Mechanics
The theory of computational mechanics proposed by Crutchfield, Shalizi, and Feldman et al. tried to formulate this kind of emergent causation in a quantitative framework Crutchfield and Young (1989). In some sense, computational mechanics can be understood as the inverse of statistical mechanics. This is because statistical mechanics derives macro-level consequences from micro-dynamics, whereas the inverse process is performed by computational mechanics, which constructs a minimal causal model from the observations of a stochastic process that can generate the observed time series.
Let us assume that the stochastic process under consideration can be represented as $\overleftrightarrow{s}$. We can divide it into two segments beginning from time step $t$: the history before $t$, denoted as $\overleftarrow{s_{t}}$, and the future after $t$, denoted as $\overrightarrow{s_{t}}$. If the process is stationary, we can remove the denotation of $t$. Thus, all possible histories $\overleftarrow{s}$ form a set, denoted as $\overleftarrow{S}$, and all futures $\overrightarrow{s}$ form a set, denoted as $\overrightarrow{S}$.
We aim to establish a model that can reconstruct and predict observed random sequences, with a higher accuracy being desirable. However, the randomness of the sequences prevents us from achieving a perfect reconstruction unless we record every randomly occurring character. This would make the model excessively long. To preserve useful information as concisely as possible, we need a coarse-grained mapping that captures the ordered structure in the random sequences, known as patterns Crutchfield (1994). We can partition $\overleftarrow{S}$ into mutually exclusive and jointly comprehensive subsets that form a set $\mathcal{R}$. Any subset $R\in\mathcal{R}$ is called a "state". We define a function from histories to states as $\eta:\overleftarrow{\mathcal{S}}\rightarrow\mathcal{R}$. Thus, $\eta$ is a method that can partition the history into mutually exclusive and jointly comprehensive subsets.
@@
What kind of state set can achieve the best balance between predictiveness and parsimony? We can introduce an equivalent relationship called causal equivalence Crutchfield (1994). Concretely, we say $\overleftarrow{s}$ and $\overleftarrow{s}^{\prime}$ are causally equivalent if and only if:
@@
This equivalent relationship can partition all the histories into equivalent classes, and they are defined as causal states. We denote all the causal states of a history $\overleftarrow{s}$ as $\epsilon(\overleftarrow{s})$, where $\epsilon:\overleftarrow{\mathcal{S}}\rightarrow 2^{\overleftarrow{\mathcal{S}}}$ is a function that can map a history $\overleftarrow{s}$ to the causal state $\epsilon(\overleftarrow{s})\in 2^{\overleftarrow{\mathcal{S}}},$ which is a subset of the histories.
For any two causal states $\mathcal{S}_{i}$ and $\mathcal{S}_{j}$, we can define causal transitions as a set of labeled probabilities $T^{(s)}_{ij}$ that represent the transition from the causal state $S_{i}$ to the causal state $S_{j}$ while emitting the symbol $s\in\mathcal{A}$ Crutchfield (1994):
@@
where $\mathcal{S}$ is the current causal state, $\mathcal{S}^{\prime}$ is its successor, and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$ denotes all the sequences of length $L=1$, in which the first symbol emitted is $s$ (we use slightly different symbols from those in Shalizi (2001)). Hence, by combining: $\mathcal{S^{\prime}}=\mathcal{S}_{j}$ and $\overrightarrow{\mathcal{S^{\prime}}}^{L=1}=s$, we ensure that all subsequent causal states will have the identical initial emitted symbol $s$. We denote the set $\{T^{(s)}_{ij}:s\in\mathcal{A}\}$ by $T$. The definition of causal transitions leads to a direct conclusion: $T^{(s)}_{ij}=\text{P}(\overleftarrow{s}s\in\mathcal{S}_{j}|\overleftarrow{s}\in\mathcal{S}_{i})=\frac{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i},\overleftarrow{s}s\in\mathcal{S}_{j})}{\text{P}(\overleftarrow{s}\in\mathcal{S}_{i})}$, where $\overleftarrow{s}s$ is read as the semi-infinite sequence obtained by concatenating the history $\overleftarrow{s}$ and the symbol $s$ Shalizi (2001).
The $\epsilon$-machine of a process is defined as the ordered pair $\left\{\epsilon,T\right\}$, where $\epsilon$ is the causal state function (which can map a state $s$ to the partition $\epsilon(s)$), and $T$ is the set of the transition matrices (the dynamics) for the states defined by $\epsilon$.
Up until now, we have defined $\epsilon$-machine, which is a pattern discovery machine where the patterns are unraveled from the set of histories. The remaining task is to show that it is, in some sense, optimal. It has the important properties of being minimally predictive, maximally statistically complex, and minimally stochastic Shalizi (2001).
We compare the causal state $\mathcal{S}$ to any state $R\in 2^{\overleftarrow{\mathcal{S}}}$ to show that for the purpose of predicting the future, the causal states do a better job, i.e., they provide more information. This property can be formulated as a mathematical theorem, called the maximal predictability theorem. It is stated as follows: if $\mathcal{S}$ is the causal state given by $\epsilon$, then for any other state $R\in 2^{\overleftarrow{\mathcal{S}}}$ and all $L\in\mathbb{Z}^{+}$, we have:
@@
where $H(\cdot|\cdot)$ is the conditional entropy, and $\overrightarrow{\mathcal{S}}^{L}$ is the $L$-length sequences for future. Thus, the inequality shows that the uncertainty of $\overrightarrow{\mathcal{S}}^{L}$ given the causal state $\mathcal{S}$ is less than all other states.
After achieving optimal predictiveness, the causal state set remains the one with minimal statistical complexity. We first introduce the notion of prescient rivals denoted as $\hat{\mathcal{R}}$, which are the states that are as predictive as the causal states; viz., for all $L\in\mathbb{Z}^{+}$, $H[{\overrightarrow{\mathcal{S}}}^{L}|\hat{\mathcal{R}}]=H[{\overrightarrow{\mathcal{S}}}^{L}|\mathcal{S}]$.
@@
Next, we show that the causal states are minimally stochastic. That is to say, compared with other competitors with the same ability to predict the future, the causal states and their transition dynamics have the least uncertainty. Then, we have the minimal stochasticity theorem, which is expressed as follows: for all prescient rivals $\hat{\mathcal{R}}$,
@@
where $\mathcal{S}^{\prime}$ and $\hat{\mathcal{R}}^{\prime}$ are the next causal state of the process and the next state, respectively. This means that the causal state and $\epsilon$-machine provide the best intrinsic determinism.
Since the causal state set is considered the best, how can we compute the causal states and $\epsilon$-machine from the observed data? The authors of Crutchfield (1994) introduced a hierarchical machine reconstruction algorithm; however, the details are not reiterated here.
Although this algorithm may not be applicable to all operational scenarios, the authors presented numerical computational results and corresponding machine reconstruction pathways for chaotic dynamics, hidden Markov models, and cellular automata as examples Crutchfield (1994); Shalizi (2001).
It is interesting to compare the theory of computational mechanics with causal emergence. Indeed, we can understand that all the histories $\overleftarrow{s}$ are micro-level states, and all the states $R\in\mathcal{R}$ are macro-states. Thus the function $\eta$ that can map a history $\overleftarrow{s}$ to a state $R$ is a possible coarse-graining strategy.
It is worth pointing out that the causal state $\epsilon(\overleftarrow{s})$ is the special state that can have at least the same predictive power as the micro-state $\overleftarrow{s}$, i.e., the full history. Therefore, $\epsilon$ is similar to the notion in the effective coarse-graining strategy in Zhang and Liu (2022) (see Section 3.6.5), and the causal transitions $T$ represent the corresponding effective macro-level dynamics. The feature of minimal stochasticity characterizes the deterministic property of the macro-dynamic. This property is characterized by effective information (EI) in causal emergence theory.
Although a clear definition and quantitative theory of emergence were not provided, the authors discussed the relationship between computational mechanics and emergence Crutchfield (1994); Shalizi and Moore (2003). In Crutchfield (1994), the authors explained that emergence can be conceptualized as a dynamic process in which a pattern acquires the ability to predictably adapt to different environments, as observed by an external observer. Additionally, they differentiated intrinsic emergence from emergence itself, as intrinsic emergence goes beyond the mere production of patterns and encompasses the formation of an embedded observer within the system through these patterns.